OpenStudy (anonymous):

Need some help with solving systems of linear equations in three variables? 3x+y+z=14 -x+2y-3z=-9 5x-y+5z=30

OpenStudy (phi):

This video shows how to do elimination http://www.khanacademy.org/math/linear-algebra/v/solving-3-equations-with-3-unknowns

OpenStudy (anonymous):

yeah i watched it and dont get it at all. I would like someone mostly just to solve the equation i posted.

OpenStudy (roadjester):

You can do that two ways, one is the old fashioned algebra 1 way where you eliminate variables (in this case, hold two constant) to solve for the third, and then substitute. The second is to use a matrix and do the Gaussian elimination using the Gauss-Jordan algorithm.

OpenStudy (mathteacher1729):

I'm guessing they want you to use Gaussian elimination? If so, start by writing the coefficients in an augmented matrix: \[ \huge \begin{bmatrix} 3 & 1 & 1 & 14 \\ -1 & 2 &-3 & -9 \\ 5 & -1 & 5 & 30 \end{bmatrix} \] You can see all of the relevant steps here: http://www.math.odu.edu/~bogacki/cgi-bin/lat.cgi by clicking "Transforming a matrix to row echelon form".

OpenStudy (anonymous):

3x+y+z=14 3(-x+2y-3z)=-9*3 add both eqation get 7y -8z=-13 3( 5x-y+5z)=3*30 5(3x+y+z)=5*14 subtract both eqa we get -8y+10z=20 again 8(7y-8z)=-13*8 7(-8y+10z)=7*20 add both of them 6z=36 z=6( put in -8y+10z=20 get y=5) y=5 x=1